# Stochastic Stability of ACIMs for Piecewise Expanding $C^{1+\varepsilon}$ Maps

### 04 Apr 2026 v1

@@ 4. Existence and Convergence of Invariant Densities
$\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}}\;\to\;0\quad\text{as }\delta\to 0.$
Applying this to $h\in BV_{1,1/p}$:
$\|\Pi_{\delta}h-h\|_{1}\;\leq\;\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}}\|h\|_{\mathcal{B}}\;\to\;0,$
@@
and therefore $\|h_{\delta}-h\|_{1}\to 0$. ∎
@@ 5. Remarks and Further Directions
In this paper, we established stochastic stability of absolutely continuous invariant measures for piecewise expanding $C^{1+\varepsilon}$ maps in the framework of generalized bounded variation spaces $BV_{1,1/p}$. The key ingredient is a Lasota–Yorke inequality uniform with respect to the perturbation parameter (Thm. 3.5), which controls the spectral behavior of the perturbed Frobenius–Perron operators $P_{\delta}$.
These results demonstrate that stochastic stability persists under minimal $C^{1+\varepsilon}$ regularity assumptions ($\varepsilon>0$), extending classical results [6] that typically require higher smoothness or more restrictive function spaces such as standard BV or Hölder.
@@
    Higher dimensions: piecewise expanding maps on $\mathbb{R}^{d}$ or manifolds, adapting $BV_{1,1/p}$ via Whitney jets.

### 11 Apr 2026 v2

@@ 4. Existence and Convergence of Invariant Densities
$\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}(I)}\to 0\quad\text{as }\delta\to 0.$
Applying this to $h\in BV_{1,1/p}(I)$:
$\|\Pi_{\delta}h-h\|_{1}\leq\|\Pi_{\delta}-\Pi\|_{\mathcal{B}\to L^{1}(I)}\|h\|_{\mathcal{B}}\to 0,$
@@
and therefore $\|h_{\delta}-h\|_{1}\to 0$.
It remains to observe that the invariant density produced here coincides with that of Theorem 4.1. Indeed, Keller–Liverani implies that for all sufficiently small $\delta>0$, the eigenspace of $P_{\delta}$ at eigenvalue $1$ is one-dimensional. Hence any nonnegative normalized fixed point of $P_{\delta}$ in $BV_{1,1/p}(I)$ must equal $\Pi_{\delta}h/\|\Pi_{\delta}h\|_{1}$, so the two constructions agree. ∎
Remark 2.
Topological mixing holds for $\tau\in\mathcal{T}([0,1];s,\varepsilon)$ whenever the Markov graph of the partition $\{a_{0},\ldots,a_{q}\}$ is aperiodic and strongly connected, in particular when every branch is full (i.e. $\tau([a_{i-1},a_{i}])=[0,1]$ for all $i$). See [8].
5. Example: Nonlinear Piecewise Expanding Map
We present a nonlinear example illustrating the applicability of our results beyond piecewise linear maps.
Example 5.1 (Nonlinear piecewise expanding map).
Let $\eta>0$ be such that $\eta<\frac{1}{2\pi}$, and define
$F(x):=2x+\eta\sin(2\pi x),\qquad x\in[0,1].$
Then
$F^{\prime}(x)=2+2\pi\eta\cos(2\pi x)\geq 2-2\pi\eta>1,$
so $F$ is strictly increasing on $[0,1]$. Since $F(0)=0$ and $F(1)=2$, by the intermediate value theorem there exists a unique $a\in(0,1)$ such that
$F(a)=1.$
Define $\tau:[0,1]\to[0,1]$ by
$\tau(x)=F(x)\bmod 1.$
Then $\tau$ can be written as a piecewise map with two branches:
$\tau(x)=\begin{cases}2x+\eta\sin(2\pi x),&x\in[0,a),\\ 2x+\eta\sin(2\pi x)-1,&x\in[a,1].\end{cases}$
Each branch extends to a $C^{1+\varepsilon}$ function on its closure, since $\sin(2\pi x)$ is smooth. Moreover,
$|\tau^{\prime}(x)|=|F^{\prime}(x)|\geq 2-2\pi\eta=:s>1,$
so $\tau$ is uniformly expanding. Since $F^{\prime}$ is smooth, it is in particular Hölder continuous, hence each branch is $C^{1+\varepsilon}$.
Furthermore, $\tau$ is strictly increasing on each branch. Since
$\tau(0)=0,\quad\tau(a)=1,$
the first branch maps $[0,a]$ onto $[0,1]$. Similarly,
$\tau(a)=0,\quad\tau(1)=1,$
so the second branch maps $[a,1]$ onto $[0,1]$. Thus both branches are full.
Therefore, $\tau\in\mathcal{T}([0,1];s,\varepsilon)$ and is topologically mixing. By Theorem 4.3, the invariant densities $h_{\delta}$ of the perturbed operators $P_{\delta}=Q_{\delta}P_{\tau}$ converge in $L^{1}(I)$ to the invariant density $h$ of $P_{\tau}$ as $\delta\to 0$.
This example shows that our results apply to genuinely nonlinear expanding systems. In particular, the use of the space $BV_{1,1/p}(I)$ allows one to treat maps with limited regularity beyond the classical bounded variation setting.
@@ 6. Remarks and Further Directions
In this paper, we established stochastic stability of absolutely continuous invariant measures for piecewise expanding $C^{1+\varepsilon}$ maps in the framework of generalized bounded variation spaces $BV_{1,1/p}(I)$. The key ingredient is a Lasota–Yorke inequality uniform with respect to the perturbation parameter (Thm. 3.6), which controls the spectral behavior of the perturbed Frobenius–Perron operators $P_{\delta}$.
These results demonstrate that stochastic stability persists under minimal $C^{1+\varepsilon}$ regularity assumptions ($\varepsilon>0$). They extend the stochastic stability result of Góra [9]—which treated the same map class via Rychlik’s $C_{\varepsilon}$ seminorm under more restrictive expansion conditions—to the $BV_{1,1/p}(I)$ framework, and complement the spectral stability theory of [4].
@@
    Higher dimensions: piecewise expanding maps on $\mathbb{R}^{d}$ or manifolds, adapting $BV_{1,1/p}(I)$ via Whitney jets.